Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Probability: A random 3 didgit no picked...

Expert replies
by Naruto » Thu Jun 11, 2009 3:00 am
What is the probability that a 3-digit positive integer picked at random will have one or more "7" in its digits?

A. 271/900

B. 27/100

C. 7/25

D. 1/9

E. 1/10



Answer given is C

My approach was to find the probabilty that 7 doesnt appear in the 3 digit no. i.e for each digit 9 nos can be chosen out of 10 (excluding 7) so the probabilty is 9x9x9/10x10x10. i.e 729/1000.
Therfore the anwer should be 1-729/1000 which is 271/1000. since its not available in the option i took the nearest value which is 27/100. Please explain what am i doing wrong here. Thanks.
Join the discussion
Source: — Problem Solving |

by rah_pandey » Thu Jun 11, 2009 3:12 am
No of 3 digit integers are not 1000 but 900 rest being one digit or 2 digit
Since it is 3 digit no therefore 0 cannot occupy hundred place. therefore no of cases are 8*9*9=648(without 7 as a possible digit)

now 900-648=252 no of 3 digits integers that have exactly one 7
reqd probability=252/900=84/300=28/100=7/25

Answer should be C
Join the discussion

by Naruto » Thu Jun 11, 2009 4:08 am
rah_pandey wrote:No of 3 digit integers are not 1000 but 900 rest being one digit or 2 digit
Since it is 3 digit no therefore 0 cannot occupy hundred place. therefore no of cases are 8*9*9=648(without 7 as a possible digit)

now 900-648=252 no of 3 digits integers that have exactly one 7
reqd probability=252/900=84/300=28/100=7/25

Answer should be C

Thanks i feel real stupid not considering the obvious.
Join the discussion

by abhinav85 » Thu Jun 11, 2009 4:35 am
Hey Rah_pandey
Join the discussion

by abhinav85 » Thu Jun 11, 2009 4:36 am
Hey rah_pandey

"No of 3 digit integers are not 1000 but 900 rest being one digit or 2 digit"

why didnt you take 999 instead of 900???
Join the discussion

by rah_pandey » Thu Jun 11, 2009 5:44 am
Total no of integers from 0-99 =100
100-999=900 therefore 900 3 digit positive integers
Join the discussion

by Osirus@VeritasPrep » Thu Jun 11, 2009 6:38 am
I see how you guys did the problem, but can someone explain why it would be incorrect to do the problem like this:

probability that 7 is in the hundreds digit = 1/9

prob that 7 is in the tens digit= 1/10

prob that 7 is in the units digit = 1/10

1/9 + 1/10 + 1/10 = 29/90

I know its not an answer choice but why is that approach wrong?
Join the discussion

by tohellandback » Thu Jun 11, 2009 6:50 am
osirus0830 wrote:I see how you guys did the problem, but can someone explain why it would be incorrect to do the problem like this:

probability that 7 is in the hundreds digit = 1/9

prob that 7 is in the tens digit= 1/10

prob that 7 is in the units digit = 1/10

1/9 + 1/10 + 1/10 = 29/90

I know its not an answer choice but why is that approach wrong?
osirus,
how about probability that 7 is in the hundreds digit and 7 is in the tens digit,
then you caculate probability that 7 is in the tens digit and 7 is in the unit's digit, and so on....
you see where you are going..

but you could solve this like this too if it makes things easier for you:

3 digit numbers containing 7s:
X07, X17, X27... X97; X= 1,2,3,4,5,6,8,9 - total will be 10*8=80
X70, X71, X72... X76, X78, X79; total will be 9*8=72
700, 701... 199 - total of 100

total=252
prob=252/900

Hope I helped
The powers of two are bloody impolite!!
Join the discussion